OpenStudy (kayla.dunton00):

I don't know how to do this question... What are the possible number of positive, negative, and complex zeros of f(x) = -2x3 - 5x2 + 6x + 4 ? Part 2 [4 points]: Use complete sentences to explain the method used to solve this equation.

OpenStudy (karim728):

solve for x or factor?

OpenStudy (kayla.dunton00):

well actually I read it wrong it says: What are the possible number of positive, negative, and complex zeros of f(x) = -2x3 - 5x2 + 6x + 4 ? Use complete sentences to explain the method used to solve this equation.

satellite73 (satellite73):

"possible" right, not actual

OpenStudy (kayla.dunton00):

yes

satellite73 (satellite73):

\[f(x) = -2x^3 - 5x^2 + 6x + 4 \] for the possible positive zeros, you are looking for the change in sign of the coefficients do you know what that means?

OpenStudy (kayla.dunton00):

yes, so there would be 1 change right?

satellite73 (satellite73):

exactly that means there must be one positive zero

satellite73 (satellite73):

now for the negative ones, look for the number of changes of sign of \(f(-x)\)

satellite73 (satellite73):

first off of course you need \(f(-x)\) do you know what that is?

OpenStudy (kayla.dunton00):

yes it would be changing all the signs opposite right? so there would still be only 1 change?

satellite73 (satellite73):

no it does not change all the signs you have to replace \(x\) by \(-x\) and see what you get

satellite73 (satellite73):

the sign of the coefficients of the odd powers will change, but the coefficients of the even powers will remain the same, as will the constant

satellite73 (satellite73):

step one is \[f(-x)=-2(-x)^2-5(-x)^2+6(-x)+4\]

satellite73 (satellite73):

ok typo there, step one is \[f(-x)=-2(-x)^3-5(-x)^2+6(-x)+4\]

OpenStudy (kayla.dunton00):

so it would be 2x^3+5x^2-6x+4

satellite73 (satellite73):

no

satellite73 (satellite73):

the \(2x^3\) is right so is the \(-6x\) and of course the \(+4\)

satellite73 (satellite73):

but \((-x)^2=x^2\) so \(-5(-x)^2=-5x^2\)

OpenStudy (kayla.dunton00):

oh i forgot to square the x's

satellite73 (satellite73):

so what do you get, and how many changes of sign does it have?

OpenStudy (kayla.dunton00):

you'd have 2x^3-5x^2-6x+4. there would be 2 changes

satellite73 (satellite73):

yes so either 2 negative zeros, or no negative zeros (you count down by twos)

satellite73 (satellite73):

in fact this one has two negative zeros and one positive one

OpenStudy (kayla.dunton00):

so there are no complex zeros?

satellite73 (satellite73):

not for this one of curse there can be at most 3 total since it is a poly of degree 3 choices would be a) one positive two complex or b) one positive, two negative

OpenStudy (kayla.dunton00):

okay

OpenStudy (kayla.dunton00):

Thank you so much!

satellite73 (satellite73):

yw